Find the value of log3 log3 log9 729 + 3 log√3 3.

Grade:11

3 Answers

Sher Mohammad IIT Delhi
10 years ago
9^3=729
log9(729)=3 3*logv3((v3)^2)=3*2=6
log3(3)=1
log3(1)=0
log3log3log9729+3logv33=6

Pradeeep Anand
10 years ago
plot of above function: the function f(x) = x-2+1=x-1 for x>2 = -x+2+1=-x+3 for x<2 } )f(x)dx=)x-1 (between 2 to x(x>2))=(1/2)x^2-x =)-x+3(between x to 2(x<2))=(1/2)(-x^2)+3x
naveen
10 years ago
6